OpenStudy (anonymous):

Is the following problem simplified correctly? (4x+5)(10x+1)=40x^2+4x+15x+5 or 40x^2+19x+20

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[(4x+5)(10x+1)=4x\times 5x+1\times 4x+5\times 10x +5\times 1\]

OpenStudy (anonymous):

you get \[20x^2+4x+50x+5\] \[20x^2+54x+5\]